### v1

# Benchmark Meta2 Evaluation - How to get it right

Abstract
Recent advancements in Language Models (LMs) have catalyzed the creation of multiple benchmarks, designed to assess these models’ general capabilities.
A crucial task, however, is assessing the validity of the benchmarks themselves.
This is most commonly done via Benchmark Agreement Testing (BAT), where new benchmarks are validated against established ones using some agreement metric (e.g., rank correlation).
Despite the crucial role of BAT for benchmark builders and consumers,
there are no standardized procedures for such agreement testing.
This deficiency can lead to invalid conclusions, fostering mistrust in benchmarks and upending the ability to properly choose the appropriate benchmark to use.
By analyzing over 40 prominent benchmarks, we demonstrate how some overlooked methodological choices can significantly influence BAT results, potentially undermining the validity of conclusions.
To address these inconsistencies, we propose a set of best practices for BAT and demonstrate how utilizing these methodologies greatly improves BAT robustness and validity.
[…]
Our findings underscore the necessity for standardized BAT, ensuring the robustness and validity of benchmark evaluations in the evolving landscape of language model research.
[…]
Leaderboard: https://huggingface.co/spaces/per/BenchBench
1 Introduction
As Language Models (LMs) increasingly excel across a broad range of tasks, new benchmarks – often measuring similar abilities – are constantly proposed. This deluge in benchmarks highlights the role of Benchmark Agreement Testing (BAT); namely, validating some benchmarks using other benchmarks by measuring an agreement metric (e.g., Pearson correlation) over their models’ scores (Liu2021DoQA).
BAT is often used to validate that a new proposed benchmark measures what it was designed to measure. The expectations from this measurement depend on the benchmark’s goal; demonstrating high agreement can serve to show that a new benchmark captures similar model abilities to established benchmarks (Lei2023S3EvalAS; Viswanathan2023Prompt2ModelGD; Chang2023DoLM; Li2024TreeEvalBE; Prabhu2024LifelongBE; He2024UltraEvalAL). High agreement can also validate that an efficient version of a benchmark (e.g., requiring less compute or labeling) measures the same thing as the original benchmark (Perlitz2023EfficientB; Polo2024tinyBenchmarksEL; Prabhu2024LifelongBE; Vivek2023AnchorPB). In contrast, if a benchmark aims to test a unique trait – one that is not properly covered by existing benchmarks – BAT will be used to demonstrate the disagreement of such benchmarks with existing ones (Yuan2024PRobELMPR; waldis2024holmes). The above goals are relevant both for benchmark creators and for benchmark consumers. Creators will typically use BAT to validate the properties of their new benchmark; benchmark consumers might use it to choose which existing benchmark they want to use.
However, despite the wide application of BAT in recent years, there is a glaring absence of common methodology. Specifically, the significance of several BAT methodological decisions is currently overlooked, undermining the validity of any conclusions made.
In this work, we aim to bring order and consistency into the practice of BAT. Analyzing more than of the most common benchmarks (§2), spanning over models, we show the critical impact of several methodological decisions in BAT, effectively altering the conclusions that researchers will draw from their analyses (§3).
We focus on three such critical choices: selecting the reference benchmark (§3.1), the models included in the test (§3.2), as well as the correlation metrics and their interpretation (§3.3). For example, as seen in Figure 1, choosing a different subset of models produces substantially different correlation scores, leading to different conclusions about benchmark agreement. The figure exemplifies that two benchmarks can (and often do) show high agreement over a wide range of models, while agreement over a few top-ranked models remains low.
[…]
To sum up, our contributions are as follows:
[leftmargin=14pt, itemsep=0mm]
1.
We perform a large-scale analysis of benchmark agreement, highlighting the impact of several crucial methodological decisions (§3).
2.
We propose guidelines for reliable and standardized BAT (§4) and demonstrate their impact.
3.
[…]
4.
[…]
2 Setup
For our analysis, we use over benchmarks, with their results cutoff at Jan 2024. The benchmarks we used include: AGI Eval (Zhong2023AGIEvalAH), Alpaca (v2) (alpaca_eval), and its length-adjusted version (Dubois2024LengthControlledAA), HuggingFace OpenLLM Leaderboard (open-llm-leaderboard), MMLU (Hendrycks2020MeasuringMM), MAGI (MAGIbenchmark), Chatbot-Arena and MTBench (Zheng2023JudgingLW), Big Bench Hard (suzgun2022challenging). HumanEval (Chen2021EvaluatingLL) ARC (clark2018think), HellaSwag (zellers2019hellaswag), TruthfulQA (lin2022truthfulqa), Winogrande (sakaguchi2019winogrande), GSM8k (cobbe2021traininggsm8k). EQ-Bench (v2) (paech2023eqbench), ArenaHard (arenahard2024) and OpenCompass (2023opencompass). For a wider survey of benchmarks used, see App. 9.1.
Our analysis focuses on evaluating agreement between two benchmarks – a reference benchmark (established and commonly acceptable) and a target benchmark (the one we assess, e.g., a new benchmark). Specifically, agreement is calculated as the correlation over the models ranks (using Kendall (Kendall1938ANM)) or scores (using Pearson (PearsonVIINO)).
We note that an inherent constraint in BAT is the size of the model intersection between the benchmarks (i.e., models appearing in both benchmarks). Benchmarks lacking a sufficiently large set of intersecting models (for this work, we chose ), cannot be reliably used for BAT.
| Recommendations | BAT Variance |
[…]
Reduction | ||||||||
[…]
3 BAT Methodological Decisions: An Analysis
When conducting BAT, researchers face a multitude of decisions: which reference benchmarks to compare against, which models to select for comparison, which metrics to use, how to define "agreement" between benchmarks, and so on.
In the absence of clear guidelines, benchmark creators often make arbitrary choices, without clear justification or consistency across different studies.
In this section, we demonstrate how such arbitrary choices hinder the validity of BAT conclusions. Next, we highlight how commonly reported BAT results can foster false expectations among benchmark consumers.
3.1 The Choice of Reference Benchmark Matters
Finding a reference benchmark for BAT is a non-trivial task. One needs to find a well-established benchmark, whose data is readily available, and which exhibits a large enough overlap with the models already evaluated in the target benchmark. Due to the above difficulty, BAT is commonly done against one or two reference benchmarks (Yuan2024PRobELMPR). Benchmarks can be divided into groups according to their measured abilities – for example, holistic benchmarks that aim to measure some loosely-defined construct of overall model quality, such as BigBench (srivastava2023beyond), benchmarks measuring coding abilities (Chen2021EvaluatingLL), math benchmarks (cobbe2021traininggsm8k) etc. Thus, when selecting a reference benchmark, there is often a somewhat arbitrary choice between several possible benchmarks which are all seemingly appropriate.
Figure 2 illustrates the variability stemming from such arbitrary choices: for each target benchmark, different reference benchmarks produce wildly varying agreement scores. For example, Alpaca V2 (second row from above) demonstrates a wide range of agreement levels with other benchmarks, spanning from a mediocre agreement of with MT-bench to a high agreement of with LMSys Arena, even though both of these reference benchmarks are considered to measure similar abilities. This variability calls into question the validity of conclusions based on applying BAT using a single reference benchmark.
To address this issue, we advocate using an aggregated reference benchmark that consolidates multiple benchmarks based on the mean-win-rate; see more on this in §4.
3.2 The Choice of Models Matters
In performing BAT, one measures some agreement metric over the scores of a group of models overlapping between the target and reference benchmark. Typically, authors arbitrarily pick some small set of models for their analysis. However, as we detail below, both the quantity and the properties of the selected models should be taken into account when drawing conclusions from BAT.
The Number of Compared Models Matters
Figure 5 demonstrates the relationship between the number of models and the variability of BAT results. It shows that with a small amount of models, BAT results can get highly unreliable, with a standard deviation approaching . For instance, in our analysis we found that the Kendall-tau correlation between LMSysArena and MT-Bench can range from approximately 0.65 to 0.99, depending on the particular number of models chosen. Thus, we see that the common practice of using a small number of models for BAT may jeopardise the validity of conclusions.
Granularity Matters
Performing BAT produces a score that indicates high or low agreement. However, the meaning of this score will differ depending on the models included in the analysis. For example, as seen in Figure 1, for a given pair of benchmarks, the agreement obtained over similarly strong models will generally be lower than over a set of models of varying qualities.
To quantify this phenomenon, we investigate benchmark agreement where the subset of models selected is not completely random, but is constrained to sets of models that are adjacent in rank (e.g., models 3-7)111Note that the sets of adjacent models were not selected from a specific rank location (e.g., Top, Bottom, Middle) but were randomly selected from the full range. For an analysis of such location-dependent sets, see App 9.2.. Adjacent models have more similar performance. Thus, their score differences and ranking may be less stable, resulting in lower correlation scores. In Figure 3, we show that indeed, for a given number of models, the correlation score when considering adjacent models is lower than that of randomly sampled models, with a stronger effect as the number of models in the subset decreases.
This discrepancy emphasizes the importance of using varying granularities when reporting BAT scores. This would enable managing the expectations of benchmark consumers, who may expect and desire a specific level of granularity (e.g., getting the very best models right, or discriminating between strong and weak models).
3.3 The Choice of Correlation Metric (and Threshold) Matters
BAT is the process of measuring correlations of model scores (or ranks) between two benchmarks. Once a correlation score is obtained, this score is commonly interpreted based on how it compares to some threshold; surpassing the threshold means the agreement is considered "high", while falling below it means the agreement is "low".
Currently, there are no consistent standards for the types and thresholds of correlation metrics. For instance, Liu2021DoQA utilized both rank and score correlations, setting a uniform threshold of for both, whereas sun2023validity exclusively employed rank correlation and opted for a distinct threshold of .
To improve our understanding on the significance of these choices, we analyse the relationship between rank (Kendall-tau) and score (Pearson) correlation metrics. In Figure 5 we present correlation scores between different pairs of benchmarks with varying model subsets. We observe a strong linear relationship () between the two correlation functions, indicating that they exhibit similar behavior in measuring agreement. However, the figure also shows a consistent score difference of approximately between the two metrics, indicating a potential flaw in the current practice of applying the same threshold regardless of the metric chosen. This underscores the necessity for a data-driven approach – comparative in nature – to interpret correlation scores; see §4 for more details.
4 BAT Best Practices
Use an Aggregate Reference Benchmark
The choice of reference benchmark can hinder the validity of BAT conclusions (§3.1). To circumvent this issue, we propose to combine the results from all benchmarks appropriate for the goal of the BAT (choosing benchmarks that measures similar or dissimilar abilities) into an aggregate reference benchmark, by averaging their model win-rates. For example, when using BAT to validate some efficient holistic benchmark, the reference benchmark should be the aggregate of all available holistic benchmarks. By combining results from a group of benchmarks, the aggregate benchmark can provide both a more stable and robust basis for comparison. Notably, since the aggregate benchmark captures the distribution of relevant results, it constitutes a better measure of the underlying construct represented by the group, called in the literature convergent validity (carlson2012understanding).
Measuring the effect of such methodology, in Table 1, we compare the standard deviation of BAT correlation results when using arbitrary reference benchmarks (first line) to that when using the aggregate, it shows that the standard deviation of the correlation drops with our recommendation by more that %.
Use More Models and Sample Them Randomly
BAT based on a small set of models has a large variance (§3.2). We recommend using at least 10 models when performing BAT, preferably more, to ensure more reliable results. These models should represent the entire spectrum of available models, including diverse sizes, architectures, and training methods. Aiming for a random selection ensures equal representation and minimizes bias. Table 1 shows that using this methodology to select models decreases BAT variance by more than %.
Report Multiple Granularities
Benchmark agreement varies significantly with the range of qualities of the models used in BAT (§3.2), see Figure 1. To manage the expectations of benchmark consumers, who may expect and desire a specific level of granularity (e.g., getting the very best models right, or discriminating between strong and weak models) we recommend reporting agreement scores at multiple resolutions (e.g., 5/10/20 contiguous models, averaging across groups when more models were sampled). This approach provides a more nuanced view on benchmark agreement, highlighting critical distinctions that might otherwise be missed (e.g. the top 3 models are almost never in agreement across benchmarks).
Use a Data-driven Threshold
Using predetermined thresholds to interpret correlation scores may be misleading, as the meaning of "high" or "low" agreement inherently has a relative aspect to it. For instance, consider the correlation scores obtained when using different model granularities. As seen in Figure 3, benchmarks produce very different correlation scores depending on the model subset chosen. Thus two benchmarks that are in low agreement can show high correlation scores in some configurations.
This difficulty calls for the use of a data-driven approach for interpreting agreement scores. Thus, for a given configuration and reference benchmark(s), we compile a distribution of agreement scores of various benchmarks against this reference. Then, we compare the correlation score obtained for the target benchmark with this distribution. Specifically, we calculate the Z-score of the target-reference correlation score relative to this distribution. A target benchmarks that produce a Z-score above are considered in agreement, whereas below it, they are not. As this method incorporates a natural distribution of benchmark agreement scores, it harnesses the natural mean agreement while also incorporating the nature of the distribution, meaning that the notion of agreement becomes relative to the population of existing benchmarks. As more benchmarks are added the test will more accurately reflect the agreement according to the current natural distribution of benchmarks.
Follow The Above Rules!
[…]
Making the case for our above recommendations, Table 1 demonstrates the significant gains obtained when using our methodological choices to perform BAT. It shows not only that the different recommendations each have an impact on variance, but also that their effect can be combined to achieve a substantially lower variance point – reducing the standard deviation by %, and thereby delivering far more robust BAT results.
[…]
The workflow of using the package is as follows:
[leftmargin=14pt, itemsep=0mm]
1.
A user inputs their BAT configuration, including the desired group of reference benchmarks.
2.
[…]
recommends a set of models for evaluation on the target benchmark.
3.
The user inputs their benchmark results for the recommended models.
4.
[…]
produces a full BAT report.
[…]
6 BAT uses in Related Work
While some examples were given in the text, we elaborate on a handful of works employing BAT.
Some works survey and analyze a field by utilizing BAT techniques. Liu2021DoQA check agreement across many QA datasets and conclude that since agreement is high, there is no need for more QA datasets. sun2023validity use correlations to show that Compositionality Benchmarks do not agree amongst themselves. They used Kendall-Tau and set as the high agreement threshold. Other works performed general efficient evaluation research and utilized BAT (Prabhu2024LifelongBE; Perlitz2023EfficientB; Polo2024tinyBenchmarksEL; Viswanathan2023Prompt2ModelGD). All of these works performed a thoughtful evaluation and large (reliable) rank correlation over all the models in the benchmarks. However, they did not consider the high correlations achieved in such settings (§3.2).
Other work relies on BAT to compare to a specific benchmark. Feng2024SampleEfficientHE automatically sample a small set of instructions as an efficient LLM benchmark, reducing human labor significantly. They show this still agrees with existing benchmarks. Similarly, Lei2023S3EvalAS and Viswanathan2023Prompt2ModelGD both propose a synthetic benchmark as a proxy and show good agreement with the original benchmark, although they differ in their methodology. Chang2023DoLM propose two benchmarks and use agreement to show that they capture the same phenomenon, and Mizrahi2023StateOW test agreement within the same benchmarks using different prompts. Li2024TreeEvalBE validate a new benchmark with 6 models of 3 sizes 7B,13B,33B with agreement alpaca(v2 )(alpaca_eval).
Last, Yuan2024PRobELMPR and (waldis2024holmes) show divergent validity by comparing their benchmark to established ones, showing low BAT scores.
7 Discussion and Conclusions
[…]
In this paper, our focus was on the methodological issues when performing BAT. We did not deal with questions regarding when BAT should be used, and how conclusions from BAT should be interpreted. Next, we describe several such open questions.
What do we make of high agreement?
It is not trivial how one should treat two benchmarks that are in high agreement with each other. If one is more convenient to run (e.g., doesn’t require costly metrics), then from a practical perspective, a user can simply choose it over the more expensive one. However, practitioners and researchers must not confuse high agreement with the notion that the benchmarks actually measure the exact same qualities. Among other things, this could lead to an erroneous view that new benchmarks are no longer needed, impeding new benchmark development. The community must also discriminate between correlations of model abilities (strong models are strong at many tasks) and correlations of the benchmarks themselves (the benchmarks actually measure the same qualities).
What do we make of low agreement?
Reliability concerns the consistency of benchmark results. In this paper, we accept the benchmark scores as presented and focus on their benchmark validity, which assesses whether benchmarks accurately measure what they purport to evaluate. However, this ignores the reliability issues within the benchmarks, which place an upper bound on the level of benchmark agreement. If, for instance, a benchmark cannot reliably differentiate between its top-3 models, then naturally we do not expect to see agreement over the top-3 models with other benchmarks. Looking forward, methodological improvements in BAT must include incorporating reliability measures, allowing to decouple disagreements from low reliability.
How do we use BAT to retire benchmarks?
Another point concerns the role of BAT for benchmark retirement, i.e., at what point do we decide that an old benchmark is no longer relevant and should be discarded. Currently the issue of retirement is viewed mainly from the perspective of saturation, where the community stops using benchmarks on which all new models succeed. However, another reason to retire benchmarks may be that the mixture of abilities models are expected to possess has shifted over time. In this scenario, BAT can reveal that a certain benchmark is no longer viable.
8 Limitations
We note that finding low agreement may mean two things, while both are negative, they should be addressed or interpreted differently. One option is that the benchmark measures something different from what it is supposed to and is hence not valid. That is the more common interpretation and calls for changes. Another option might be that the benchmark is just not reliable, intuitively its ranking is unstable and did not converge. In such cases, even the same benchmark may not agree with itself given small changes (subsets, seeds etc.), this usually calls for evaluating on more examples (choshen-etal-2024-navigating). There is a positive note to the same story, if a benchmark already shows a strong BAT in fine-grained evaluation (e.g., 5 models close to each other), it also means that it is quite reliable.
Sometimes BAT is not needed. BAT gives a way to validate a benchmark by an external source of authority. However, other methods or other sources for authority (e.g., being masterfully crafted by experts) might give stronger signals. Especially in the case of new and unique signals that can mostly show they are different, but not that they are valid for their own unique purpose.
In general, BAT needs a reference benchmark, or ideally multiple ones giving diverse measurements of the same construct. Still, choosing the right reference benchmarks might be tricky, and the results might be sensitive to this choice.
9 Appendices
9.1 Benchmarks used
The AGI Eval (Zhong2023AGIEvalAH) benchmark assesses models on human-level cognition and problem-solving tasks, which tests the real-world applicability of model outputs. Similarly, Alpaca (v2) (alpaca_eval) and its length-adjusted version (Dubois2024LengthControlledAA) focus on a model’s ability to follow complex instructions with the latter specifically addressing biases associated with output length.
HumanEval (Chen2021EvaluatingLL) presents code generation challenges, evaluating the syntactic correctness and logical soundness of model-generated code. Alongside, the HuggingFace OpenLLM Leaderboard (open-llm-leaderboard) employs the Eleuther AI Evaluation Harness (eval-harness) to test models on several key benchmarks such as ARC (clark2018think), HellaSwag (zellers2019hellaswag), MMLU (hendrycks2021measuring), TruthfulQA (lin2022truthfulqa), Winogrande (sakaguchi2021winogrande), and GSM8k (cobbe2021gsm8k). EQ-Bench (v2) (paech2023eqbench), measures the emotional intelligence of models, essential for applications that involve nuanced human interactions.
The MAGI (MAGIbenchmark) benchmark integrates challenging elements from MMLU and AGIEval to test complex reasoning and problem-solving capabilities of models. It is particularly effective in highlighting subtle performance differences among top-tier models. MMLU (Hendrycks2020MeasuringMM) assesses both general and specialized knowledge across various domains, providing a broad evaluation spectrum.
Further, benchmarks like Chatbot-Arena and MTBench (Zheng2023JudgingLW) focus on multi-turn conversation abilities, crucial for applications in customer service and virtual assistance. Lastly, Big Bench Hard (suzgun2022challenging) challenges models with complex text understanding and generation, pushing the limits of what natural language processing technologies can achieve. It is worth noting, that the HELM benchmark (liang2023holistic) was excluded from our analysis because there were few overlapping models with the other benchmarks.
9.2 Model Tier
Building on the importance of model proximity, another crucial factor in benchmark agreement is the tier of models being assessed. Current BAT practices often treat benchmarks as a uniform slab, disregarding the variations across different tiers of model performance. However, agreement might not be uniform across these tiers, and understanding this variance can provide deeper insights into benchmark reliability and model performance.
In Figure 7, we show that model tier significantly impacts benchmark agreement. Bottom-tier models exhibit higher agreement among themselves, with Kendall correlation coefficients just below 0.5. In contrast, middle-tier models show low agreement (coefficients below 0.2), and top-tier models demonstrate low to medium agreement (around 0.3).
One potential explanation for this phenomenon is the (lack of) reliability of the benchmark, as discussed in the introduction and literature (Perlitz2023EfficientB). Figure LABEL:fig:top_middle_bottom_std highlights that the standard deviation of scores buttom-ranked models is significantly higher than the rest. This might mean that there is some effect the goes beyond granularity or density, with older models being easier to differentiate (and gaining higher correlations to the models). However middle and top ranked models do not show such a trend (even when taking into account that middle granularity is higher as top models are still joining the game), which means that no strong conclusion should be made excluding older models, switching benchmarks frequently or similar actions, at most, old models may be left out of BAT, but other effects seem more pressing.
9.3 Benchmark used for visualizations
The benchmarks we used include: AGI Eval (Zhong2023AGIEvalAH), Alpaca (v2) (alpaca_eval), and its length-adjusted version (Dubois2024LengthControlledAA), HuggingFace OpenLLM Leaderboard (open-llm-leaderboard), MMLU (Hendrycks2020MeasuringMM), Chatbot-Arena and MTBench (Zheng2023JudgingLW), Big Bench Hard (suzgun2022challenging). ARC (clark2018think), HellaSwag (zellers2019hellaswag), TruthfulQA (lin2022truthfulqa), Winogrande (sakaguchi2019winogrande), EQ-Bench (v2) (paech2023eqbench).
NeurIPS Paper Checklist
1.
Claims
x
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
x
[…]
x
Justification: Sections 3,4,5
x
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
x
Question: Does the paper discuss the limitations of the work performed by the authors?
x
[…]
x
Justification: Limitations section
x
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory Assumptions and Proofs
x
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental Result Reproducibility
x
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
x
[…]
x
Justification: section 2
x
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
x
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
x
[…]
x
Justification: Upon acceptance, company does not allow to submit annonymized
x
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental Setting/Details
x
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
x
[…]
x
Justification: Section 2
x
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment Statistical Significance
x
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
x
[…]
x
Justification: Yes
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments Compute Resources
x
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code Of Ethics
x
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
x
[…]
x
Justification: Yes
x
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader Impacts
x
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
x
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
x
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
x
[…]
x
Justification: Section 2 and Appendix
x
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New Assets
x
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and Research with Human Subjects
x
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
x
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
x
[…]
x
Justification: NA
x
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.